OpenStudy (anonymous):

What are the intercepts of 3x-9y=-27 ?

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

-9

jigglypuff314 (jigglypuff314):

x-intercept is when y = 0 y-intercept is when x = 0 so to find the x intercept plug in 0 for y to get 3x - 9(0) = -27 and solve :) then to find the y intercept plug in 0 for x to get 3(0) - 9y = -27 and solve :)

OpenStudy (the_fizicx99):

The x and y intercepts of just the y intercept?

jigglypuff314 (jigglypuff314):

@tHe_FiZiCx99 it appears to be a linear equation and is asking for plural "intercepts"

OpenStudy (anonymous):

so combine them ? it'd equal zero right 0-0=-27

jigglypuff314 (jigglypuff314):

not quite :) don't combine them, solve them separately 3x - 9(0) = -27 and solve by itself 3(0) - 9y = -27 and solve by itself

OpenStudy (anonymous):

3x- 9(0)=-27 3x- 0= -27 2(0)- 9y=-27 0 - 9y=-27

jigglypuff314 (jigglypuff314):

yep :) and you can continue 3x- 0= -27 simplifies to 3x = -27 then divide both sides by 3 0 - 9y=-27 simplifies to -9y = -27 then divide both sides by negative 9

OpenStudy (anonymous):

-9/3=-3 ? -9/-9=1

jigglypuff314 (jigglypuff314):

mmm 3x = -27 then divide both sides by 3 means /3 /3 so x = -27/3 = ? -9y = -27 then divide both sides by negative 9 means /-9 /-9 so y = 27/9 = ? both are still separate

OpenStudy (anonymous):

x=-9 y=3

jigglypuff314 (jigglypuff314):

yep! :)

OpenStudy (anonymous):

there has to be 2 sets of paired number ?

jigglypuff314 (jigglypuff314):

so like x-intercept is when y = 0 makes x = -9 ordered pairs are like (x, y) so you get (-9, 0) then you would do the same for the y = 3 part y-intercept is when x = 0 makes y = 3 ordered pairs are like (x, y) so you get (0, 3)

OpenStudy (anonymous):

Ohh ok so what about this one? What are the intercepts of x + 5y =10

jigglypuff314 (jigglypuff314):

same idea :) to get x-intercept plug in y = 0 so x + 5(0) = 10 and simplify

OpenStudy (anonymous):

x+0= 10

jigglypuff314 (jigglypuff314):

and x+0 = just x so you can take the 0 out to simplify so you get x = ?

OpenStudy (anonymous):

1?

jigglypuff314 (jigglypuff314):

errr 10 says 10 so x = 10 I was talking about the x+0 ^ part before. so since x-intercept is when y = 0 and we got that x = 10 and ordered pairs are like (x, y) how about you give it a try? :3

OpenStudy (anonymous):

so it'd be 10,0 0,2 ?

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

Hey check this one out . Rachel can buy licorice sticks for $0.075 and cerry candies for $0.50 and has a budget of $25 . If her expense are represented by the equation 0.75x+0.5y=25, where x is the number of licorice sticks and y is the number of cherry candies, how many cherry candies can she buy if she wishes to spend the enitre $25 on cherry candies? \[Note: x \ge 0, y \ge 0\], and x and y take only integer values.)

OpenStudy (anonymous):

my guess is that you divide 25 by 0.5 right ?

jigglypuff314 (jigglypuff314):

yep! :)

OpenStudy (anonymous):

ythanks can you give me a medal ? (:

jigglypuff314 (jigglypuff314):

lol ;)